Simons Center for Geometry and Physics Arts Program

Report to Board of Trustees

October 25, 2013

1. Administration. Since January 2012 the SCGP Arts Program has been without a Director. Exhibitions have been curated by free-lance personnel, coordinated by the Advisory Committee and administered by Elyce Winters. This was the situation reported in May 2012 and it endures. Our attempts to recruit for an interim half-time position for this academic year never got off the ground, reportedly because of bottlenecks in the campus Human Resources administration. We are now following the Provost’s suggestion of searching for a shared Visiting Scholar position with the Art Department. Together with John Lutterbie, the Art Chair, we have drafted a job description (in brief, a 5-year appointment at the Visiting Assistant Professor level, duties split 50-50 between Art and the SCGP), which has been submitted to Human Resources for approval. We hope for a speedy resolution so we can begin actively recruiting for a September 2014 starting date.

2. Meanwhile, the SCGP Arts Program Advisory Committee has continued, with the same membership, their role in soliciting and screening proposals for exhibitions, concerts and other arts-oriented activities. Current members are Stephanie Dinkins (Studio Art), George Hart (free-lance computer scientist and sculptor), John Lutterbie (Theatre Arts and Art) Tony Phillips (Math, Chair) and Daniel Weymouth (Music).

3. Recent exhibitions and events include

• “Nowhere differentiable,” a show of fractals in traditional African Art (Stephanie Dinkins, curator), with a lecture by the Ethno-mathematician Ron Eglash from RPI. (February 2013)

• “Paths: charting, navigating and bridging” (Alastair Noble, curator) with a lecture by the artist W. Bradford Paley. (April 2013)

• “Dance your Dissertation” a talk with video Cutting Sequences on the Double Pentagon by Diana Davis. Her video won the Physics and Math category of the 2012 “Dance Your PhD” competition, sponsored by Science and TED. This event was paired with a concert by Re’ut Ben-Ze’ev (mezzo-soprano), Andrey Tchekmazov (cello) and Leon Livshin (piano). (July 23)

• “Science and Metaphor: At Play,” a workshop on science and poetry led by the Canadian poet Alice Major (September 18). This event was organized by Prof. Alfred Goldhaber (Physics) and was coordinated with a talk by Ms. Major at the SBU Humanities Institute.

• “Euclid’s Kiss,” Geometric Sculptures of George Hart, currently on display.

4. Concerts. Weekly recitals were organized during last summer’s SCGP Graduate Summer School and the Simons Workshop. We plan to continue them next summer, with the help of pianist Leon Livshin and jazz trombonist Ray Anderson (Music), who plan programs, recruit fellow artists and also perform.

Respectfully submitted,

Tony Phillips, Arts Program Advisory Committee Chair